[Date Prev][Date Next][Thread Prev][Thread Next][Search] [Main Index] [Thread Index] [HEASARC Mailing List Archives]

CGRO biWeekly Status Report

Compton Observatory Science Report #156 Friday, April 29, 1994
Chris Shrader, Compton Observatory Science Support Center

Questions or comments can be sent to the Compton SSC.
Phone: 301/286-8434
e-mail:   NSI_DECnet: GROSSC::SHRADER 
          Internet: shrader@grossc.gsfc.nasa.gov

                         Spacecraft Status

The Observatory, scientific instruments and all spacecraft
subsystems continue to function nearly flawlessly.

                    Science Support Center News

The 238 Guest Investigator proposals received for Cycle-4 are being
processed. Preparation is underway for the peer-review which will
take place in late June.

A CD Rom containing the Phase-1 EGRET sky survey data is being
prepared for distribution at the Minneapolis AAS meeting.

                        Instrument Reports


EGRET operations were normal this biweekly period. The average
percentage of possible data recovered remains about 80% or above.
Interactions with guest investigators remains at a good level. All
EGRET data through period 210 has been transferred to the Compton
Observatory Science Support Center in accordance with the planned
delivery schedule.

As promised two weeks ago, several papers on EGRET results were
presented at the APS meeting, including ones on galactic sources,
bursts, and AGNs. As expected, the paper on the February 17, 1994
burst, wherein high energy gamma rays were seen for about an hour,
and one gamma ray had an energy of 25 GeV, received a substantial
amount of attention. This result combined with earlier ones
obviously significantly reduces the number of possible explanations
of these bursts.

Several EGRET publications have appeared recently including one on
observations of normal galaxies. The microsecond burst paper and
one on pulsars have been accepted for publication in the
Astrophysical Journal.


OSSE operations are normal.

In viewing period 324 (19-26 Apr), the Z-axis target was the
galactic plane near (l,b) = (+20,0) (Key project), and the X-axis
target was NGC 7469 (Guest Investigator R. Cameron).  When neither
target was above the horizon, the target was 4C 15.05.  The Sun was
near the OSSE scan plane, and the slewing response to BATSE solar
flare triggers was enabled.  OSSE responded to a flare trigger at
TJD 9464.01.

In viewing period 325 (26 Apr - 10 May), the Z-axis targets are GT
0236+610 (Guest Investigators M. Coe, M. Tavani) and NGC 1275, and
the X-axis targets are QSO 0438-436, PSR 0540-693, and SN 1987A (PI
team). The Sun is not available on the scan plane.

Data from viewing period 207 were delivered to the Compton GRO
Science Support Center archive this week.  The targets during this
period were IC 4329A and PSR 1822-09.


BATSE contribution to the GROSSC bi-weekly report:

The following report was included in IAU Circular  5977:

     GRO J2014+34                                                   
     M. H. Finger, Universities Space Research Association; M. 
     Stollberg and G. N. Pendleton, University of Alabama,
     Huntsville; R. B. Wilson, Marshall Space Flight Center, NASA;
     and D. Chakrabarty, J. Chiu, and T. A. Prince, California
     Institute of Technology, report for the Compton Gamma-Ray
     Observatory BATSE team: "Pulsed hard x-ray emission is being
     seen from a previously uncatalogued object in the
     Cygnus-Vulpecula region. The current best-fit source position
     is R.A. = 20h14m, Decl. = +34o.5 (equinox 2000.0), with an
     estimated 68-percent-confidence region of radius 8 deg. Work
     is continuing to verify this position estimate. The emission
     from the pulsar was first detectable on Apr. 6 and reached
     peak intensity on Apr. 16. The source is detected at energies
     from 20 up to at least 70 keV. The barycentric pulse period on
     Apr. 14.0 UT was 18.7021 +/- 0.0001 s, with the period
     increasing at a rate of (6.0 +/- 0.7) x 10E-9. The pulse
     profile averaged over Apr. 14-16 has a main peak with a FWHM
     of about 0.5 in phase, with a 0.15-HWHM 'shoulder' present on
     the falling edge. The 20- to 40-keV pulsed flux on Apr. 16 was
     about 50 mCrab, with the spectrum well fitted by an optically
     thin thermal bremsstrahlung model F(E) = (A/E)exp(-E/kT), with
     temperature kT = 23 +/- 6 keV and F(40 keV) = 1.88 +/- 0.24 x
     10E-4 photon cmE-2 sE-1 keVE-1."      
The following has been submitted to as for inclusion as an IAU

     HARD X-RAY PULSAR GRO J1948+32 (= GRO J2014+34) D. Chakrabarty
     and T. A. Prince, Caltech; M. H. Finger, Universities Space
     Research Association; R. B. Wilson, NASA Marshall Space Flight
     Center; and G. N. Pendleton, University of Alabama,
     Huntsville, report for the Compton GRO/BATSE team: "Further
     observations of the recently discovered 18.7 s transient hard
     X-ray pulsar in Cygnus (originally designated GRO J2014+34;
     see IAUC 5977) have yielded a more precise position. The
     estimated 90-percent-confidence error box is approximately a
     2.5 deg x 5.5 deg quadrilateral, with corners RA = 19h58m, Dec
     = +29 00'; 19h50m, +34 00'; 19h34m, +35 00'; 19h46m, +28 00'
     (equinox 2000.0). In accordance with the substantially
     improved position, we have redesignated the source as GRO
     J1948+32. The pulsed emission was first detectable on Apr. 6
     UT and continues to be visible as of Apr. 28. Optical and soft
     X-ray observations of the error box are strongly encouraged in
     order to further constrain the X-ray position and identify the
     accreting companion.

As of April 27, BATSE has detected 973 cosmic gamma-ray bursts out
of a total of 2846 on-board triggers in 1085 days of operation.
There have been 730 triggers due to solar flares with emission
above 60 keV.


The COMPTEL operations group reports that the instrument is
performing well and continues scheduled observations.

Ongoing scientific analysis of the available incoming flight data
is proceeding routinely at the various COMPTEL sites (i.e., there
is nothing terribly extraordinary to report this week). 

Communications continue with the staff of the Science Support
Center regarding the technical details of the installation of the
COMPASS analysis package at the SSC.

The most exciting upcoming event: preparations are well underway
for the GRO/COMPTEL solar eclipse party at the University of New
Hampshire on Tuesday, May 10. Weather gods permitting, festivities
will begin promptly at 13:42:43 EDT, when the total annular phase
begins, and last until 13:44:58 EDT, when totality ends.  All
friends of GRO and COMPTEL are cordially invited (BYO filters).